# FedHERO: A Federated Learning Approach for Node Classification Task on Heterophilic Graphs

### v1

@@ Appendix A Experiment Setups
For graph partitioning, we employ well-established algorithms commonly used in federated graph learning. Specifically, we utilize the METIS algorithm Karypis & Kumar (1997) and the Louvain algorithm Blondel et al. (2008). With METIS, we ensure an even distribution of nodes across subsets, with each subset representing a client’s data in a federated learning context. Using Louvain, we follow the methodology of Baek et al. Baek et al. (2023) and Zhang et al. Zhang et al. (2021). We initially partition the graph and select subgraphs containing a minimum of 50 nodes. The smaller subgraphs are then randomly merged into these larger ones to guarantee adequate training data. For example, if Louvain results in $20$ subgraphs but only $5$ have over $50$ nodes, we consolidate the remaining $15$ subgraphs into these $5$, effectively simulating a federated learning scenario with $5$ clients.
We further provide the description of real-world datasets. The syn-cora dataset Zhu et al. (2020) generates nodes and edges based on the expected homophily ratio within the graph, with feature vectors of nodes in each class derived by sampling from the corresponding class in the Cora dataset Yang et al. (2016). The ieee-fraud dataset comprises transaction records for products, with edges determined by the proximity of transaction occurrences. The objective is to predict fraudulent transactions Howard et al. (2019). Meanwhile, the credit dataset has 30,000 nodes, each representing an individual, and their connections are determined by the similarity of their spending and payment patterns Yeh & Lien (2009). The objective is to predict whether an individual will default on their credit card payment or not.
@@
We follow the previous work Mao et al. (2023) to compute the homophily ratios of both the original graphs and the averaged homophily ratios of partitioned graphs. As in Table 7, the partitioned graphs exhibit comparable homophily ratios when compared to the original graphs. It means that the partitioning process does not compromise the heterophilic property of the graphs.
@@ Appendix B Neighbor node distributions of clients in two datasets
In this section, we also give an example in two real-world datasets: Squirrel and Chameleon Rozemberczki et al. (2021). In these datasets, nodes correspond to web pages, and edges represent hyperlinks between them. To simulate the FL scenario, we apply the community detection algorithm METIS Karypis & Kumar (1997) to partition these graphs, and variations in the structures of the resulting subgraphs also arise. As illustrated in Figure 6, we compare the neighbor distributions of two clients in two datasets, respectively. (The neighbor distributions across all clients can refer to Figure 7.) Each bar illustrates, for a specific class of nodes, the ratio of edges connecting them to other types of nodes of one type to the total number of edges serving this class node as an endpoint. For instance, in Figure 6 (a), the orange bar with the x-axis ’5-1’ signifies the proportion of edges linking class 5 nodes to class 1 nodes in relation to the overall number of edges originating from class 5 nodes. The results indicate distinct neighbor distributions among different clients. For instance, in the Chameleon dataset, almost 70% of edges originating from class 1 nodes are intra-class connections in client 1, implying class 1 nodes tend to establish hyperlinks with similar web pages. Conversely, less than 10% of the neighbors of class 1 nodes are nodes in the same class in client 2, indicating a preference for establishing hyperlinks with diverse web pages. To safeguard data privacy, the structural details of subgraphs cannot be directly shared between clients. Consequently, this leads to distinct message-passing patterns of information transfer between these subgraphs. Traditional FL methods like FedAvg McMahan et al. (2017) simply aggregate GNN models, potentially hindering the global model’s ability to account for the aggregation discrepancies among various subgraphs effectively.
@@ C.1 Overlapping subgraphs scenario
Despite the commonality of the non-overlapping node assumption, some research investigates scenarios with overlapping nodes. We have followed the data partitioning approach from Baek et al. (2023) in our experiments, focusing on the Flickr and Actor datasets. The results, as illustrated in Table 8, show FedHERO consistently outperforming most baselines across these datasets. This indicates its robust capability to manage overlapping scenarios effectively. We theorize that overlapping subgraphs provide client subgraphs with additional connections, hinting at the potential for developing FL algorithms specifically designed to cater to overlapping environments.
@@ C.2 Latent graph generation
We also examine the influence of another prevalent latent graph generation method Wu et al. (2020a); Elinas et al. (2020); Zhao et al. (2023). This method treats each edge in the latent graph as a Bernoulli random variable, resulting in edge distributions represented by a product of $N\times N$ independent Bernoulli variables.
@@ C.3 Federated graph learning baselines
  • GCFL Xie et al. (2021): It incorporates a gradient sequence-based clustering mechanism using dynamic time warping. This mechanism enables multiple data owners with non-IID structured graphs and diverse features to collaboratively train robust graph neural networks.
  • FedStar Tan et al. (2023): It relies on structural embeddings and a feature-structure decoupled graph neural network to address non-IID data challenges through shared structural knowledge. Within the FL framework, it exclusively exchanges the structure encoder. This allows clients to collectively learn global structural knowledge while maintaining personalized learning of node representations based on their features.
  • FedLit Xie et al. (2023): It tackles the variability in link characteristics within graphs, where ostensibly uniform links may convey varying degrees of similarity or significance. It dynamically identifies latent link types through an EM-based clustering algorithm during federated learning and adapts message-passing strategies using multiple convolution channels tailored to different link types.
  • FedPub Baek et al. (2023): It leverages functional embeddings of the local GNNs with random graphs as inputs to calculate similarities between them. These similarities are then employed for weighted averaging during server-side aggregation. Additionally, the method learns a personalized sparse mask at each client, allowing for the selection and update of only the subset of aggregated parameters relevant to the specific subgraph.
  • FedSage Zhang et al. (2021): It introduces NeighGen, an innovative missing neighbor generator integrated with federated training procedures. This endeavor is geared towards achieving a universal node classification model within a distributed subgraph framework, eliminating the need for direct data sharing.
@@ Appendix D Influence of Hyperparameters in FedHERO
The architecture of the GNN models consists of a feature projection layer, two graph convolutional layers, and a classifier layer. We employ Adam optimizer Kingma & Ba (2014) with a learning rate of 0.005. We apply a consistent hyperparameter setting in our main results. Specifically, we set $\alpha=0.2$, $\mu=\lambda=0.1$, $k=20$, and $N_{H}=4$. The training process involves 200 communication rounds, with each local training epoch lasting 1 iteration. The nodes are randomly distributed into five groups. For each experiment, one group is chosen as the test set, while the remaining groups are utilized for training and validation. We perform five experiments on each dataset and report the models’ average performance. For the ieee-fraud and credit datasets, we present the mean test Area Under the Curve (AUC), accounting for the imbalanced labels. For the remaining datasets, we use mean test accuracy across clients for evaluation.
@@ D.2 Influence of Smoothing Hyperparameters $\lambda$ and $\mu$
No distinct trend related to the intrinsic nature of the data was observed for $\lambda$ and $\mu$. However, it is noteworthy that removing regularization on structures (i.e., setting $\lambda$ or $\mu$ to 00) resulted in more pronounced performance degradation. In fact, the regularization loss plays a constructive role in offering guidance for structure learning, as highlighted in prior work Zhao et al. (2023). Conversely, increasing $\lambda$ or $\mu$ to a certain value proved beneficial in enhancing the performance of FedHERO. However, further increases had a negative impact on the model’s focus on the classification task, leading to a decline in performance.
@@ D.3 Influence of Aggregation Hyperparameters $\alpha$
For the Squirrel, Chameleon, and Actor datasets, optimal results were observed at $\alpha=0.2$, while $\alpha=0.4$ exhibited the best performance for the Flickr dataset, as illustrated in Figure 9(c). This variation in optimal $\alpha$ values is attributed to the homophily ratios of the datasets. According to Mao et al.Mao et al. (2023) and Kim et al.Kim & Oh (2022), the homophily ratios for Squirrel, Chameleon, and Actor are $0.22$, $0.25$, and $0.22$, respectively. In contrast, Flickr has a higher homophily ratio of $0.32$. Given Flickr’s relatively elevated homophily ratio, employing a larger $\alpha$ to increase the proportion of local models proves beneficial, effectively leveraging the structural information within the subgraph. Conversely, for datasets with lower homophily ratios, superior performance is achieved by placing more emphasis on the homophily properties introduced by the global structure learner. In summary, we posit that, for graph data with a larger homophily ratio (indicating lower heterophily), higher values of $\alpha$ can be advantageous.
@@ Appendix E Comparative Analysis with FedStar
This subsection offers a detailed comparison between FedHERO and FedStar due to their similarity. We observe that the design of FedHERO shares similarities with the concept in FedStar Tan et al. (2023), which involves decoupling the GNN into feature and structure encoders. However, FedHERO distinguishes itself from FedStar in both purpose and the utilization of structural information. In FedStar, structural information is integrated into features to capture additional domain-invariant characteristics Tan et al. (2023). Conversely, FedHERO is tailored for scenarios where graphs in clients originate from the same domain but possess distinct node neighbor distributions. Its objective is to learn generally applicable message-passing patterns that are conducive to different heterophilic graphs and can be shared across clients. While FedStar introduces degree-based and random walk-based structure embedding techniques to extract structural information for each node, these embeddings might prove simplistic for the problem outlined in the Introduction. For instance, if one node has three neighbor nodes with the same class and another node has three neighbor nodes with different classes, the degree-based method would yield identical embeddings for both nodes. Additionally, in situations where two graphs share the same adjacency matrix but feature different node class distributions, the random walk-based method might not reach its potential. In contrast, the structure learning model in FedHERO is designed to generate a desired graph structure, bypassing the challenges encountered by the structure embedding method in FedStar. This is corroborated by the experimental results outlined in the Experiment Section.

### v2

@@ Appendix A Experiment Setups
For graph partitioning, we employ well-established algorithms commonly used in federated graph learning. Specifically, we utilize the METIS algorithm (Karypis & Kumar, 1997) and the Louvain algorithm (Blondel et al., 2008). With METIS, we ensure an even distribution of nodes across subsets, with each subset representing a client’s data in a federated learning context. Using Louvain, we follow the methodology of Baek et al. (2023) and (Zhang et al., 2021). We initially partition the graph and select subgraphs containing a minimum of 50 nodes. The smaller subgraphs are then randomly merged into these larger ones to guarantee adequate training data. For example, if Louvain results in $20$ subgraphs but only $5$ have over $50$ nodes, we consolidate the remaining $15$ subgraphs into these $5$, effectively simulating a federated learning scenario with $5$ clients.
We further provide the description of real-world datasets. The syn-cora dataset (Zhu et al., 2020) generates nodes and edges based on the expected homophily ratio within the graph, with feature vectors of nodes in each class derived by sampling from the corresponding class in the Cora dataset (Yang et al., 2016). The ieee-fraud dataset comprises transaction records for products, with edges determined by the proximity of transaction occurrences. The objective is to predict fraudulent transactions (Howard et al., 2019). Meanwhile, the credit dataset has 30,000 nodes, each representing an individual, and their connections are determined by the similarity of their spending and payment patterns (Yeh & Lien, 2009). The objective is to predict whether an individual will default on their credit card payment or not.
@@
We follow the previous work (Mao et al., 2023) to compute the homophily ratios of both the original graphs and the averaged homophily ratios of partitioned graphs. As in Table 7, the partitioned graphs exhibit comparable homophily ratios when compared to the original graphs. It means that the partitioning process does not compromise the heterophilic property of the graphs.
@@ Appendix B Neighbor node distributions of clients in two datasets
In this section, we also give an example in two real-world datasets: Squirrel and Chameleon (Rozemberczki et al., 2021). In these datasets, nodes correspond to web pages, and edges represent hyperlinks between them. To simulate the FL scenario, we apply the community detection algorithm METIS (Karypis & Kumar, 1997) to partition these graphs, and variations in the structures of the resulting subgraphs also arise. As illustrated in Figure 6, we compare the neighbor distributions of two clients in two datasets, respectively. (The neighbor distributions across all clients can refer to Figure 7.) Each bar illustrates, for a specific class of nodes, the ratio of edges connecting them to other types of nodes of one type to the total number of edges serving this class node as an endpoint. For instance, in Figure 6 (a), the orange bar with the x-axis ’5-1’ signifies the proportion of edges linking class 5 nodes to class 1 nodes in relation to the overall number of edges originating from class 5 nodes. The results indicate distinct neighbor distributions among different clients. For instance, in the Chameleon dataset, almost 70% of edges originating from class 1 nodes are intra-class connections in client 1, implying class 1 nodes tend to establish hyperlinks with similar web pages. Conversely, less than 10% of the neighbors of class 1 nodes are nodes in the same class in client 2, indicating a preference for establishing hyperlinks with diverse web pages. To safeguard data privacy, the structural details of subgraphs cannot be directly shared between clients. Consequently, this leads to distinct message-passing patterns of information transfer between these subgraphs. Traditional FL methods like FedAvg (McMahan et al., 2017) simply aggregate GNN models, potentially hindering the global model’s ability to account for the aggregation discrepancies among various subgraphs effectively.
@@ C.1 Overlapping subgraphs scenario
Despite the commonality of the non-overlapping node assumption, some research investigates scenarios with overlapping nodes. We have followed the data partitioning approach from (Baek et al., 2023) in our experiments, focusing on the Flickr and Actor datasets. The results, as illustrated in Table 8, show FedHERO consistently outperforming most baselines across these datasets. This indicates its robust capability to manage overlapping scenarios effectively. We theorize that overlapping subgraphs provide client subgraphs with additional connections, hinting at the potential for developing FL algorithms specifically designed to cater to overlapping environments.
@@ C.2 Latent graph generation
We also examine the influence of another prevalent latent graph generation method (Wu et al., 2020a; Elinas et al., 2020; Zhao et al., 2023). This method treats each edge in the latent graph as a Bernoulli random variable, resulting in edge distributions represented by a product of $N\times N$ independent Bernoulli variables.
@@ C.3 Federated graph learning baselines
  • GCFL (Xie et al., 2021): It incorporates a gradient sequence-based clustering mechanism using dynamic time warping. This mechanism enables multiple data owners with non-IID structured graphs and diverse features to collaboratively train robust graph neural networks.
  • FedStar (Tan et al., 2023): It relies on structural embeddings and a feature-structure decoupled graph neural network to address non-IID data challenges through shared structural knowledge. Within the FL framework, it exclusively exchanges the structure encoder. This allows clients to collectively learn global structural knowledge while maintaining personalized learning of node representations based on their features.
  • FedLit (Xie et al., 2023): It tackles the variability in link characteristics within graphs, where ostensibly uniform links may convey varying degrees of similarity or significance. It dynamically identifies latent link types through an EM-based clustering algorithm during federated learning and adapts message-passing strategies using multiple convolution channels tailored to different link types.
  • FedPub (Baek et al., 2023): It leverages functional embeddings of the local GNNs with random graphs as inputs to calculate similarities between them. These similarities are then employed for weighted averaging during server-side aggregation. Additionally, the method learns a personalized sparse mask at each client, allowing for the selection and update of only the subset of aggregated parameters relevant to the specific subgraph.
  • FedSage (Zhang et al., 2021): It introduces NeighGen, an innovative missing neighbor generator integrated with federated training procedures. This endeavor is geared towards achieving a universal node classification model within a distributed subgraph framework, eliminating the need for direct data sharing.
@@ Appendix D Influence of Hyperparameters in FedHERO
The architecture of the GNN models consists of a feature projection layer, two graph convolutional layers, and a classifier layer. We employ Adam optimizer (Kingma & Ba, 2014) with a learning rate of 0.005. We apply a consistent hyperparameter setting in our main results. Specifically, we set $\alpha=0.2$, $\mu=\lambda=0.1$, $k=20$, and $N_{H}=4$. The training process involves 200 communication rounds, with each local training epoch lasting 1 iteration. The nodes are randomly distributed into five groups. For each experiment, one group is chosen as the test set, while the remaining groups are utilized for training and validation. We perform five experiments on each dataset and report the models’ average performance. For the ieee-fraud and credit datasets, we present the mean test Area Under the Curve (AUC), accounting for the imbalanced labels. For the remaining datasets, we use mean test accuracy across clients for evaluation.
@@ D.2 Influence of Smoothing Hyperparameters $\lambda$ and $\mu$
No distinct trend related to the intrinsic nature of the data was observed for $\lambda$ and $\mu$. However, it is noteworthy that removing regularization on structures (i.e., setting $\lambda$ or $\mu$ to 00) resulted in more pronounced performance degradation. In fact, the regularization loss plays a constructive role in offering guidance for structure learning, as highlighted in prior work (Zhao et al., 2023). Conversely, increasing $\lambda$ or $\mu$ to a certain value proved beneficial in enhancing the performance of FedHERO. However, further increases had a negative impact on the model’s focus on the classification task, leading to a decline in performance.
@@ D.3 Influence of Aggregation Hyperparameters $\alpha$
For the Squirrel, Chameleon, and Actor datasets, optimal results were observed at $\alpha=0.2$, while $\alpha=0.4$ exhibited the best performance for the Flickr dataset, as illustrated in Figure 9(c). This variation in optimal $\alpha$ values is attributed to the homophily ratios of the datasets. According to Mao et al. (2023) and Kim & Oh (2022), the homophily ratios for Squirrel, Chameleon, and Actor are $0.22$, $0.25$, and $0.22$, respectively. In contrast, Flickr has a higher homophily ratio of $0.32$. Given Flickr’s relatively elevated homophily ratio, employing a larger $\alpha$ to increase the proportion of local models proves beneficial, effectively leveraging the structural information within the subgraph. Conversely, for datasets with lower homophily ratios, superior performance is achieved by placing more emphasis on the homophily properties introduced by the global structure learner. In summary, we posit that, for graph data with a larger homophily ratio (indicating lower heterophily), higher values of $\alpha$ can be advantageous.
@@ Appendix E Comparative Analysis with FedStar
This subsection offers a detailed comparison between FedHERO and FedStar due to their similarity. We observe that the design of FedHERO shares similarities with the concept in FedStar (Tan et al., 2023), which involves decoupling the GNN into feature and structure encoders. However, FedHERO distinguishes itself from FedStar in both purpose and the utilization of structural information. In FedStar, structural information is integrated into features to capture additional domain-invariant characteristics (Tan et al., 2023). Conversely, FedHERO is tailored for scenarios where graphs in clients originate from the same domain but possess distinct node neighbor distributions. Its objective is to learn generally applicable message-passing patterns that are conducive to different heterophilic graphs and can be shared across clients. While FedStar introduces degree-based and random walk-based structure embedding techniques to extract structural information for each node, these embeddings might prove simplistic for the problem outlined in the Introduction. For instance, if one node has three neighbor nodes with the same class and another node has three neighbor nodes with different classes, the degree-based method would yield identical embeddings for both nodes. Additionally, in situations where two graphs share the same adjacency matrix but feature different node class distributions, the random walk-based method might not reach its potential. In contrast, the structure learning model in FedHERO is designed to generate a desired graph structure, bypassing the challenges encountered by the structure embedding method in FedStar. This is corroborated by the experimental results outlined in the Experiment Section.